OpenStudy (anonymous):

First make a substitution, then integrate by parts to evaluate the integral: integral of x*ln(1+x) dx

ganeshie8 (ganeshie8):

u = x^2/2 v = ln(1+x)

OpenStudy (anonymous):

isn't that just going straight into integration by parts?

ganeshie8 (ganeshie8):

oh, by substitution do they really mean a 'substitution method' ?

OpenStudy (anonymous):

text is asking to make a u-sub, then integrate that new integrand. It's confusing me - is it just a fancy way of saying integrate by parts anyway?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

now u making me doubt... wid that u-sub first... lol

OpenStudy (anonymous):

it can be solved that way it's all over the net, but using sub method first is nowhere to be found

ganeshie8 (ganeshie8):

i dont see any advantage in u-sub in any way -.-

OpenStudy (anonymous):

lol i don't either, I'll ask professor thanks anyway

ganeshie8 (ganeshie8):

okie